OpenStudy (anonymous):

Eight inches of heavy wet snow are equivalent to two inches of rain. Estimate the water content in 9 inches of wet snow

OpenStudy (goformit100):

Form The Equation First.

OpenStudy (anonymous):

this is where I am stuck not sure how I would think 8 = 2

OpenStudy (anonymous):

I know this isn't right

OpenStudy (nincompoop):

\[\frac{ 8 }{ 2 }:\frac{ 9 }{ x }\]

OpenStudy (anonymous):

I should have known that I am solving for x

OpenStudy (nincompoop):

it doesn't have to be x, it could be anything. it's just that x is commonly used to pertain to an unknown

OpenStudy (anonymous):

its 8/2 + 9/x if i am not mistaken though

OpenStudy (anonymous):

it would \[\frac{ 9 }{ 4 }\]

OpenStudy (anonymous):

but wouldn't I divide

OpenStudy (nincompoop):

try cross multiplication \[\frac{ 8 }{ 2 }:\frac{ 9 }{ x }=18=8x\] now solve for x

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\]

OpenStudy (nincompoop):

no... you were almost there. you got \[\frac{ 9 }{ 4 }\] because you reduced your \[\frac{ 8 }{ 2 }=\frac{ 4 }{ 1 }=4\]

OpenStudy (anonymous):

I have to type an integer or a decimal I'm working on homework thank you for your help

OpenStudy (anonymous):

so its 4 not 1/2

OpenStudy (nincompoop):

okay let us go back to what I typed earlier \[\frac{ 8 }{ 2 }:\frac{ 9 }{ x }=18=8x\] divide both sides by 8 to get the value of x

OpenStudy (anonymous):

\[\frac{ 4 }{ 8 }\]

OpenStudy (nincompoop):

\[\frac{ 18 }{ 8 } = x\]

OpenStudy (anonymous):

damn it

OpenStudy (nincompoop):

try it again with another problem

OpenStudy (anonymous):

Im so lost

OpenStudy (nincompoop):

I am sorry

OpenStudy (nincompoop):

do you know what it means to cross multiply?

OpenStudy (anonymous):

I do

OpenStudy (anonymous):

I don't know why that wasn't clicking

OpenStudy (nincompoop):

\[\frac{ a }{ b }:\frac{ c }{ d }\rightarrow ad = bc\]

OpenStudy (anonymous):

now would I divide \[\frac{ 18 }{ 8 }\]

OpenStudy (anonymous):

I do have to enter a decimal or an integer

OpenStudy (anonymous):

when divided I got 2.25

OpenStudy (nincompoop):

YES :) a decimal or an integer would be fine. unless you are being asked for a specific one. you are dividing by 8 because you had 8x and you are looking for the value of x \[\frac{ 8x }{ 8 }=\frac{ 18 }{ 8 }\] look on the left-hand side of the equation 8 ÷ 8 = 1 1x is the same as x right?

OpenStudy (nincompoop):

but I think you meant FRACTION or DECIMAL

OpenStudy (anonymous):

oh ok awesome thank you for getting me through this

OpenStudy (nincompoop):

\[\frac{ 18 }{ 8 }=2\frac{ 2 }{ 8 }=2\frac{ 1 }{ 4 }=2.25\]

OpenStudy (anonymous):

2.25